# DPO: Dual-Perturbation Optimization for Test-time Adaptation in 3D Object Detection

### v1

@@ 4.5. Efficiency Analysis
To assess the efficiency of DPO, we conducted a comparative analysis with selected TTA baselines focusing on adapting speed (i.e., seconds per frame) and GPU memory usage, as illustrated in Figure 6. CoTTA is identified as the most efficient TTA baseline for 3D object detection, demonstrating rapid adaptation capabilities (under 0.15 seconds per frame). Conversely, other baselines, notably SAR, required significantly more adaptation time and yet underperformed, achieving $\text{AP}_{\text{BEV}}$ of less than 66%. Despite a slightly longer processing time, DPO markedly surpassed all TTA baselines, showcasing its superior performance. In terms of GPU memory consumption, CoTTA reported moderate usage, whereas MemCLR exceeded 18,000 MiB but fell short in performance. The proposed DPO, in contrast, not only required less GPU memory than both MemCLR and CoTTA but also achieved dominating adaptation performance, highlighting the efficiency and effectiveness of our method.
@@ 5. Conclusion
In this work, we present a novel framework for Test-Time Adaptation in 3D Object Detection (TTA-3OD) aimed at adapting 3D detection models to new unlabeled scenes during the inference phase with a single pass. Our approach incorporates worst-case perturbations at both model and input levels to enhance robustness and generalization, thereby enabling 3D detectors to stably adapt to any test scenes with corruptions. We employ reliable Hungarian matching for trustworthy pseudo-label selection, with an early cutoff to avoid computation burden and error accumulation. Our proposed approach achieves state-of-the-art performance on three real-world transfer tasks. Beyond point- and voxel-representation-based 3D detectors used in this paper, our future work will further validate multimodal detectors with different input modalities, such as BEVfusion (Liu et al., 2023) to verify shifts across modalities.
@@ B.2. Additional Implementation Details
Tent and SAR (Wang et al., 2021; Niu et al., 2023) utilize entropy minimization to optimize the batchnorm layers during test time. Therefore, we calculate the entropy loss by summing the classification logits for all proposals at the first detection stage. CoTTA (Wang et al., 2022) follows a mean-teacher framework. Although a broad range of data augmentations is typically required to generalize the model to various corruptions, our empirical evidence from test-time adaptation for 3D object detection (TTA-3OD) suggests that most augmentations do not improve—and may even impair—performance. The sole exception is random world scaling. As a result, we adopt random world scaling as our primary strategy, in accordance with (Luo et al., 2021), applying strong scaling (0.9 to 1.1) and weak scaling (0.95 to 1.05), respectively. Regarding pseudo-labeling, we directly apply strategies tailored for 3D object detection from (Yang et al., 2021, 2022) to enhance self-training in CoTTA. Similar to CoTTA, we adopt the same augmentation strategy for MemCLR (VS et al., 2023), which was originally tailored for image-based 2D object detection, and extend it to 3D detection scenarios. This involves reading and writing pooled region of interest (RoI) features extracted during the second detection stage and computing the memory-based contrastive loss. We apply all hyperparameters from the original paper by default. Besides, The proposed DPO is a pseudo-labeling-based self-training approach for TTA-3OD. We leverage the self-training paradigm and augmentation strategies from prior works (Yang et al., 2021, 2022; Luo et al., 2021). The complete configuration files and implementation code are included in this supplementary material.

### v2

@@ 4.5. Efficiency Analysis
To assess the efficiency of DPO, we conducted a comparative analysis for adaptation speed (i.e., seconds per frame) and GPU memory usage, as illustrated in Figure 6. CoTTA is identified as the most efficient TTA baseline for 3D object detection, demonstrating rapid adaptation capabilities (under 0.15 seconds per frame). Conversely, other baselines, notably SAR, required significantly more adaptation time and yet underperformed, achieving $\text{AP}_{\text{BEV}}$ of less than 66%. Despite a slightly longer processing time, DPO markedly surpassed all TTA baselines, showcasing its superior performance. In terms of GPU memory consumption, CoTTA reported moderate usage, whereas MemCLR exceeded 18,000 MiB but fell short in performance. The proposed DPO, in contrast, not only required less GPU memory than both MemCLR and CoTTA but also achieved dominating adaptation performance, highlighting the efficiency and effectiveness of our method.
@@ 5. Conclusion
In this work, we present a novel framework for Test-Time Adaptation in 3D Object Detection (TTA-3OD) aimed at adapting detectors to new unlabeled scenes with a single pass. Our approach incorporates worst-case perturbations at both model and input levels to enhance robustness and generalization, thereby enabling 3D detectors to stably adapt to any test scenes with corruptions. We employ reliable Hungarian matching for trustworthy pseudo-label selection, with an early cutoff to avoid computation burden and error accumulation. Beyond point- and voxel-representation-based 3D detectors used in this paper, our future work will further validate multimodal detectors with different input modalities, such as BEVfusion (Liu et al., 2023) to verify shifts across modalities.
Acknowledgements.
This research is partially supported by the Australian Research Council (DE240100105, DP240101814, DP230101196, DP230101753).
@@ B.2. Additional Implementation Details
Tent and SAR (Wang et al., 2021; Niu et al., 2023) utilize entropy minimization to optimize the batchnorm layers during test time. Therefore, we calculate the entropy loss by summing the classification logits for all proposals at the first detection stage. CoTTA (Wang et al., 2022a) follows a mean-teacher framework. Although a broad range of data augmentations is typically required to generalize the model to various corruptions, our empirical evidence from test-time adaptation for 3D object detection (TTA-3OD) suggests that most augmentations do not improve—and may even impair—performance. The sole exception is random world scaling. As a result, we adopt random world scaling as our primary strategy, in accordance with (Luo et al., 2021), applying strong scaling (0.9 to 1.1) and weak scaling (0.95 to 1.05), respectively. Regarding pseudo-labeling, we directly apply strategies tailored for 3D object detection from (Yang et al., 2021, 2022) to enhance self-training in CoTTA. Similar to CoTTA, we adopt the same augmentation strategy for MemCLR (VS et al., 2023), which was originally tailored for image-based 2D object detection, and extend it to 3D detection scenarios. This involves reading and writing pooled region of interest (RoI) features extracted during the second detection stage and computing the memory-based contrastive loss. We apply all hyperparameters from the original paper by default. Besides, The proposed DPO is a pseudo-labeling-based self-training approach for TTA-3OD. We leverage the self-training paradigm and augmentation strategies from prior works (Yang et al., 2021, 2022; Luo et al., 2021).